OpenStudy (prupel85):

if the complex number x=3+bi and lxl^2=13 which is a possible value of b? A:2 B:4 C:9 D:10

hartnn (hartnn):

|x| is called the modulus of a complex number. If x = a+bi \(|x| = \sqrt{a^2+b^2}\) so what will be |x|^2 when x = 3+bi ?

OpenStudy (prupel85):

i have no clue lol

hartnn (hartnn):

so, when x = 3+bi \(|x|^2 = 3^2 + b^2\) now this is given to be 13 so 3^2+b^2 = 13 can you find 2 possible values of b from this?

OpenStudy (prupel85):

b=2?

hartnn (hartnn):

very good! :) correct.

OpenStudy (prupel85):

thanks for the help